OpenStudy (anonymous):

will give medal 6x^2+7x=0 separate the two values by commas find the solution set

OpenStudy (lastdaywork):

\[6x^2+7x=0\] On factorizing it, we get \[x(6x+7)=0\] Can you tell me the answer now ??

OpenStudy (anonymous):

i need fractions to replace the x's

OpenStudy (lastdaywork):

You can also (separately) equate each factor to zero and solve for x.

OpenStudy (anonymous):

how?

OpenStudy (lastdaywork):

\[x(6x+7)=0\] implies either x = 0 ; or 6x + 7 = 0 Do you know how to solve linear equations ?

OpenStudy (anonymous):

kind of

OpenStudy (lastdaywork):

Solve 6x + 7 = 0

OpenStudy (anonymous):

ok thanks

OpenStudy (lastdaywork):

You're welcome :)